OpenStudy (anonymous):

Explain? Solve for x: 4/3 x + 2/3 = 6

OpenStudy (schrodinger):

I would so be willing to help you out, if you didn't have that profile picture, lol. Not quite so sure about the implications of "you need a history lesson".

OpenStudy (anonymous):

So, because of where I come from and being myself, you wont help? Thanks.

OpenStudy (schrodinger):

Back. Yes, because where i'm from, your people don't like my people. Where i'm from, your people treated my people atrociously, and as if they were subhuman. I don't attach cultural ideas to where i'm from, simply because I happen to be a particular race doesn't mean I ascribe to ideas, simply because i'm from a geographical area doesn't mean i'm consciously forced to identify with that region's politics. This is nothing about where you're from, it's that you're choosing to culturally support something that is bigoted.

OpenStudy (schrodinger):

In fact, I think i'm going to fetch an administrator. I'm not so cool with symbols of slavery. That's just not cool. At all.

OpenStudy (anonymous):

I dont mean to offend anyone, I'm just a redneck, I cant help it.

OpenStudy (schrodinger):

You can. You could easily choose a different picture. You could choose an innocuous one. I mean, please.

OpenStudy (anonymous):

I will change it then, sorry it offended you.

OpenStudy (schrodinger):

Whatever. I'm just going to ignore this and walk away, but this is just a little messed up. If admins saw a profile with a nazi flag, it would be a different story. No problem. I'd be happy to help you still, but that flag makes me think of things that i'd really rather not think about.

OpenStudy (anonymous):

Just remember, I dont mean it in that way. I grew up around this flag, raised around rednecks, I diagree with the stuff it mkes younthink about, but honestly, I think it looks cool also, another reason why I love it . Hah.

OpenStudy (schrodinger):

No problem. People grow up in certain environments, it's just that that means a whole different world to some people. I'm still a little bothered, but we're all here to learn. That being said, x: 4/3 x + 2/3 = 6. In order to find the answer, let's make all of the numbers in our problem have the same denominator, which in this case would be 3. We'll need to do this to 6, so let's take six, multiply it by 3, and then put three in the denominator, but don't simplify it. So, what would you get from this?

OpenStudy (anonymous):

18/3

OpenStudy (schrodinger):

Great. Okay, so from here we can easily add fractions that have the same denominator. Given 4/3x + 2/3 = 18/3 we can subtract 2/3rds from the right side, which gives us 4/3x = 16/3. Oh, bu the way, clarification, is this \[4\div3x \]Or\[(4/3)x\]?

OpenStudy (schrodinger):

*by the way

OpenStudy (anonymous):

4/3 x + 2/3 = 6\[\frac{ 4 }{ 3 }x + \frac{ 2 }{ 3 } = 6\]Subtract 2/3 from both sides. \[\frac{ 4 }{ 3 }x = 6 - \frac{ 2 }{3 }\]Multiply both sides by 3. \[4x = 18 - 2\]\[4x = 16\]Divide both sides by 4. \[x = 16/4\] What is x?

OpenStudy (schrodinger):

^(...Or that, but i'm really not sure if you meant \[(4/3)x\], or if you meant \[\frac{ 4 }{ 3x }\]

OpenStudy (anonymous):

Yeah, it depends what you meant when you wrote the equation. (:

OpenStudy (schrodinger):

@Panda223RulerZ

OpenStudy (schrodinger):

Oh well. Well, if you have any issues, get back to either of us.

OpenStudy (anonymous):

Sorry, trying to cook XD But I understand it, thanks guys(:

OpenStudy (schrodinger):

No problem. Thanks for changing your profile picture. I didn't mean to be a bother about it.

OpenStudy (anonymous):

Its fine, sorry to offend you

OpenStudy (shadowfiend):

Folks, I just want to say, well handled. Tempers ran high for a minute there, but you peacefully discussed and resolved the issues and managed to get some learning and teaching done in the process. I was here the entire time and at no point did I feel the need to step in. Huge huge props. You are an example to everyone on this site—and, indeed, to folks all over the world.

OpenStudy (farmdawgnation):

I want to echo @shadowfiend's sentiments. You both did a superb job of resolving your differences without having the entire conversation degrade into senseless drama. Well done.